OpenStudy (anonymous):

Calculate the ratio of diffusion rates for neon and helium. Which gas differs? About how much faster? Show your steps.

OpenStudy (jfraser):

use the equation of graham's law\[\frac {R_2}{R_1} = \sqrt\frac{MM_1}{MM_2}\]

OpenStudy (chmvijay):

Diffusion of gas is inversely proportional to their respective molecular weight

OpenStudy (chmvijay):

lower the molecular weight diffuses faster and higher the molecular weight diffuses slow

OpenStudy (chmvijay):

can you make out now for neon and helium by comparing their molecular weight :)